News Flash Home
The original item was published from 12/7/2015 12:05:00 PM to 12/7/2015 12:05:23 PM.

News Flash

Community News

Posted on: November 2, 2015

[ARCHIVED] Installation Schedule For Well #3 Pipe (Updated 12/03/15)

Update 12/03/15
The well has been put back together and is now flowing to the water treatment plant. Current flow rates are slightly higher than anticipated so some minor adjustments may be made over the next 24 hours to accommodate for the additional flow.

The Village will continue flowing the well and begin easing back into normal operations. Depending upon where residents are located in town and there water usage in their particular area it may take 24-48 hours for all of the existing water to be used in the distribution system.

Minor work on closing up the exterior of the building will be completed tomorrow 12/03/15.

Update 12/03/15
Final preparations are ongoing for bringing the well back to the treatment plant and distribution. The well is anticipated to be brought back online sometime this afternoon. An updated post with additional information will be made once that occurs.

Update 12/02/15
The well has been closed back up this afternoon and is tentatively scheduled to begin going back to production tomorrow pending no other unforeseen issues. Additional information on the run time and well operations will be posted tomorrow morning.

Update 12/02/15
The Village has received confirmation that two samples have cleared testing on the well. Crews are out today re-assembling the well. Crews are scheduled will be on site Thursday or Friday to close the well building back up. Updated information regarding the well going back to distribution should be available later today.

Update 11/30/15
The Village will continue to flush the well today and will take another sample to send to the lab for testing. The first sample sent last week did test positive for traces of bacteria, which is not unusual when re-starting a well after a maintenance period. The well must pass two lab tests 24 hours apart before it can be put back to distribution. The well will continue to be flushed and samples will be sent daily to the lab until the standard requirements are met.

Update 11/25/15
The well was flushed yesterday and the first sample was taken and sent to the lab. Flushing of the well will continue today.

Update 11/24/15
The well is running and is currently flushing to the sewer while being monitored. A sample will be taken later this afternoon and sent to a lab for testing.

Update 11/23/15
Crews were on site today and hooked up the equipment needed for flushing the well. An electrician was on site and determined there was a fault in the control transformer for the well. The unit has been removed and will need to be bench tested. Unfortunately, flushing cannot proceed until the control transformer is cleared.

Update 11/21/15
The well pipe installation has been completed this afternoon and the well has been put back together. Staff will begin flushing the well on Monday morning and then take water samples to a lab for testing (labs are not open on the weekend). Results from the lab tests may take 24-72 hours, but once all samples have cleared the well will be put back into production.

Update 11/20/15
Work crews were onsite last night until after 9:00 pm pulling up the pipe and they located the problem associated with the poor electrical readings. Crews will be back onsite today reinstalling the pipe and they will continue to work this evening as weather conditions allow. Work will be scheduled to continue tomorrow (Saturday) however the start time and duration may be dependant upon the severity of the upcoming snow storm.

Update 11/19/15
Installation of the pump and pipe received a setback today. After setting the pipe for today crews onsite tested the electrical system around noon and the resulting megger reading showed only 0.51 megohms which means there is a problem somewhere. Testing of the electrical system yesterday (11/18) showed normal readings. Unfortunately to track down the issue crews will have to remove the pipe they installed today to inspect the cabling run down through the pipe.

The direct impact upon the timing and finalization of the pipe and pump installation won't be known until the problem is found. Additional updates will be posted as they become available.

Update 11/19/15
Significant progress was made on 11/18 with the installation of the well pipe. Installation will continue today with the remainder of the well pipe with the target to begin testing the well on Friday, 11/20.

Update 11/17/15
Due to rain and wind conditions work has been halted today. Work will continue pending the weather on 11/18/15.

Update 11/16/15
Installation of the pump and motor is ongoing today. Work for installing the remaining pipe is expected to be completed on Tuesday or Wednesday. The well will then be started up and undergo testing. Once testing is complete and samples come back clear (24-72 hours typically) the well will be put back into service.

Update 11/13/15
Once again due to high wind gusts, work has been postponed due to safety concerns hoisting the pipe in the high winds. Crews are anticipated to be onsite tomorrow weather permitting and work will continue into early next week.

Update: 11/12/15
Due to high winds crews are unable to perform work today by hoisting and lifting the well pipe. The overall completion schedule has been delayed by one day.

Update: 11/09/15
Site mobilization begins today at well #3. A brush cleaning of the well shaft liner will commence and then be re-televised. The well pipe will then be set following the cleaning. Work is anticipated to be completed by the end of this week, but is dependent upon the final cleaning process. Re-activation of the well may be between 11/13 to 11/18 depending upon final test results for water quality.

The Village has received an updated installation schedule from Layne-Christensen for the re-installation of the well #3 pipe.

Pump repairs for the well have been ongoing for the past several weeks. Reconstruction of the damaged bowl for the pump is expected to be completed during the middle of this week and epoxy coating of the well pipe should be completed by the end of this week.

The well rig and pump will be moved back on site the week of November 9th at which time work to clean the liner and televise the well will begin. Installation of the well pipe and testing of the pump will take place after that with the target completion date of November 13th.

Please note that unforeseen circumstances may result in timelines being shifted. Updated information will be posted when it is available.

Facebook Twitter Email